If tana=P/Q,prove that Q cos2a+P cos2a=Q
1
Expert's answer
2019-05-02T05:24:19-0400

Let's try:

if

"\text{tan}\alpha=\frac{P}{Q},"

then

"Q\text{cos}(2\alpha)+P \text{cos}(2\alpha)=\text{cos}(2\alpha)(Q+P)=Q."


First, try to check whether it is true substituting real numbers. Take "\alpha=30^\circ" for example:


"\text{tan}30^\circ=\frac{1}{\sqrt{3}},"

or "P=1,\space\space\space Q=\sqrt{3}." Now substitute these values into what we must prove:


"\text{cos}(2\cdot30^\circ)(\sqrt{3}+1)=1.366,"

and it is not equal to "Q=\sqrt{3}."

Perhaps there is a typo and instead of what we wrote we must write


"Q\text{cos}^2(\alpha)+P \text{cos}^2(\alpha)=\text{cos}^2(\alpha)(Q+P)=Q."

Check this with the same angle:


"\text{cos}^2(30^\circ)(\sqrt{3}+1)=2.049,"

which again has nothing common with "Q=\sqrt{3}=1.732."

You can try it with other combination of cosines squared or with double angles and the closest result can be obtained when


"Q\text{cos}^2(\alpha)+P \text{cos}(2\alpha)=Q:"

"\sqrt{3}\text{cos}^2(30^\circ)+P \text{cos}(2\cdot30^\circ)=1.799,"

but it is again different from Q.

This cannot be proved.